# Neural Graph Generator: Feature-Conditioned Graph Generation using Latent Diffusion Models

### v2

@@ Abstract
Graph generation has emerged as a crucial task in machine learning, with significant challenges in generating graphs that accurately reflect specific properties. Existing methods often fall short in efficiently addressing this need as they struggle with the high-dimensional complexity and varied nature of graph properties. In this paper, we introduce the Neural Graph Generator (NGG), a novel approach which utilizes conditioned latent diffusion models for graph generation. NGG demonstrates a remarkable capacity to model complex graph patterns, offering control over the graph generation process. NGG employs a variational graph autoencoder for graph compression and a diffusion process in the latent vector space, guided by vectors summarizing graph statistics. We demonstrate NGG’s versatility across various graph generation tasks, showing its capability to capture desired graph properties and generalize to unseen graphs. This work signifies a significant shift in graph generation methodologies, offering a more practical and efficient solution for generating diverse types of graphs with specific characteristics.
@@ 1 Introduction
Recent graph generative models have improved in accurately representing various properties of real-world graphs. Nonetheless, to produce graphs adhering to specific properties, it is necessary first to compile a dataset of graphs that display these characteristics before proceeding to train the model using this dataset. Thus, to generate different types of graphs, different models need to be trained, which is impractical. This work aims to fill this gap. Specifically, we develop a neural network model, so-called Neural Graph Generator (NGG), to generate graphs that exhibit specific properties. We capitalize on recent advances in latent diffusion models to perform vector-conditioned graph generation. The model applies the diffusion process not to the graph data but instead to an encoded latent representation of the graph. We produce vectors that contain a summary of the statistics of each graph (e. g., number of nodes, number of edges, clustering coefficient, etc.). Those vectors guide the denoising part of the diffusion process. We first pre-train a variational graph autoencoder which we use to map graphs into vectors and map vectors into graphs. The diffusion is performed in the space of vectors. To generate new graphs that exhibit specific properties, we sample a vector from the standard normal distribution, and we denoise this vector, while denoising is guided by the vector of graph properties. Once the noise has been removed from the vector, we feed the emerging vector into the decoder of the pre-trained variational graph autoencoder to produce a graph. The proposed model is evaluated in the task of graph generation conditioned on graph properties. Our results indicate that the NGG model can indeed generate graphs whose properties are similar to the ones that are given as input to the model. To summarize, our work makes the following contributions:
@@ 2 Related Work
For the problem of graph generation, as discussed above, most models belong to one of the following five families: (1) Auto-regressive models; (2) Variational autoencoders; (3) Generative adversarial networks; (4) Normalizing flows; and (5) Diffusion models. Auto-regressive models assume a specific node ordering and generate graphs in a sequential fashion. GraphRNN [You et al., 2018], GraphGen [Goyal et al., 2020] and GraphGen-Redux [Bacciu and Podda, 2021] are three examples of auto-regressive models. GraphRNN generates nodes and its associated edges sequentially, while the other two models generate edges and their endpoint nodes sequentially. Variational autoencoders consist of two modules. First, the encoder which maps the input data to a space that corresponds to the parameters of a Gaussian distribution. Typically, this distribution is encouraged to be similar to the standard normal distribution. Second, the decoder which maps from the latent space to the input space. In the graph domain, GraphVAE [Simonovsky and Komodakis, 2018] is a model whose encoder is an instance of a message passing graph neural network, while its decoder is a simple multi-layer perceptron. There are models that embed graphs into distributions different from Gaussian such as DGVAE [Li et al., 2020] which utilizes the Dirichlet distributions as priors on the latent variables and the latent variables represent graph cluster memberships. Variational autoencoders have also achieved success in the field of chemoinformatics [Jin et al., 2018, Samanta et al., 2020]. Similar to variational autoencoders, generative adversarial networks also consist of two main components, a generator whose objective is to generate realistic samples and a discriminator whose objective is to distinguish real samples from synthetic ones. Most models from this family typically consist of a generator that takes vectors sampled from a standard normal distribution and processes them with a multi-layer perceptron to generate a graph, and of a discriminator which is an instance of a message passing graph neural network [De Cao and Kipf, 2018, Pölsterl and Wachinger, 2021]. Architectures based on normalizing flows explicitly model a probability distribution by leveraging a method that uses the change-of-variable law of probabilities to transform a simple distribution into a complex one by a sequence of invertible and differentiable mappings. Normalizing flows can be applied to graphs by designing message passing mechanisms that are exactly reversible [Liu et al., 2019]. Such models have been applied to molecular data. For example, MoFlow [Zang and Wang, 2020] generates chemical bonds through a Glow-based model, and atoms given bonds using a graph conditional flow. Diffusion models gradually add noise to the input data in the forward diffusion process, and then learn to remove the noise in the reverse diffusion process. There exist models that add Gaussian noise to the graph’s adjacency matrix and binarize the continuous values of the output of the reverse diffusion process to produce valid graphs [Niu et al., 2020, Jo et al., 2022]. More recent models apply a discrete diffusion process that progressively adds noise by adding or removing edges and changing the nodes’ and/or edges’ types [Vignac et al., 2023].
@@ 3 Neural Graph Generator
Figure 1: Overview of the proposed architecture. The variational graph autoencoder is responsible for generating a compressed latent representation for each graph. Those representations are fed to the diffusion model which operates in that latent space. The denoising process is conditioned on a vector that contains the graph’s properties. The output of the diffusion model is passed on to the decoder which generates a graph.
@@ 4 Experimental Evaluation
With regard to the variational autoencoder, the encoder $\mathcal{E}$ consists of $2$ layers, the hidden dimension size is set equal to $64$, while the input graphs are embedded into $32$-dimensional vectors. We imposed an ordering on the nodes based on the BFS tree’s default ordering. The decoder $\mathcal{D}$ consists of $3$ layers and the hidden dimension size is set to $256$. The output of $\mathcal{D}$ is a $100^{2}$-dimensional vector and is transformed into an adjacency matrix of dimension $(100\times 100)$. Thus, the model can handle graphs of up to $100$ nodes. We also initially annotate the nodes of each graph with $10$-dimensional spectral features (from the eigenvectors associated with the $10$ smallest eigenvalues of the normalized Laplacian matrix). For the latent diffusion model, we set the number of timesteps $T$ equal to $500$, while we use a cosine schedule for adding noise at each time step. The denoising neural network $\epsilon_{\theta}$ consists of $3$ layers, the hidden dimension size is equal to $512$, while the size of the output is set to $128$. To train both models, we use the Adam optimizer and we set the learning rate to $0.001$. The batch size is set to $256$. We set the number of epochs of the variational autoencoder to $200$, while the number of epochs of the latent diffusion model to $100$. Overall, the variational autoencoder consists of $2,640,492$ parameters, while the diffusion model consists of $973,088$ parameters. Thus, in total, the NGG model consists of $3,613,580$ parameters.
@@
We also use a variant of the variational autoencoder (VGAE) as a baseline where the process is conditioned on vector of properties $\mathbf{c}$ for controlled sampling by concatenating the vector with the graph’s latent representation.
@@
Assortativity coefficient 0.30 61.32 0.11 39.10
# triangles 10,356.20 99.91 1,026.44 24.38
Avg. # triangles formed by an edge 8.85 66.52 9.44 68.32
Max. # triangles formed by an edge 539.03 83.97 49.26 16.66
Avg. local clustering coefficient 0.29 35.04 0.08 15.42
Global clustering coefficient 0.36 58.12 0.05 14.07
@@ 4.1 Within Distribution Performance
# nodes 15.48 18.76 6.68 6.66 25.70 26.05 27.30 27.54
# edges 355.79 60.12 99.09 11.31 785.20 49.83 832.26 52.90
Density 0.28 48.10 0.07 10.74 0.27 34.04 0.29 36.65
Min. degree 12.59 67.12 7.95 41.49 13.96 57.71 13.36 57.50
Max degree 4.46 7.95 2.93 4.64 23.97 35.17 26.04 38.06
Avg. degree 14.34 51.23 3.11 9.27 18.83 40.84 20.17 43.77
Assortativity coefficient 0.29 65.66 0.72 47.57 0.46 60.98 0.43 65.18
# triangles 3,913.93 83.79 1,482.23 27.21 15,101.24 74.92 15,956.96 79.93
Avg . # triangles formed by an edge 7.78 54.99 17.93 74.95 9.11 49.11 17.73 77.18
Max . # triangles formed by an edge 278.36 64.51 83.15 17.71 709.65 66.55 762.29 71.23
Avg. local clustering coefficient 0.26 26.01 0.09 15.91 0.33 40.78 0.36 44.69
Global clustering coefficient 0.27 46.16 0.07 14.44 0.32 41.89 0.34 44.34
Max k-core 11.39 52.05 2.28 9.99 16.93 42.96 17.99 45.76
# communities 2.64 27.47 1.02 12.32 1.86 22.98 1.96 23.40
Diameter 3.52 29.98 2.55 16.61 3.40 23.82 3.31 24.19
All 0.96 56.13 0.54 28.89 0.77 42.14 0.78 42.88
@@ 5 Conclusion
We introduced the Neural Graph Generator, a conditional latent diffusion model for efficient and accurate graph generation. NGG represents a significant new paradigm, adeptly generating graphs with specific, user-defined properties, a task that has long challenged traditional models. We demonstrated the ability of NGG to generalize in graphs than those in the training set, and the ability to handle missing values in the condition vector. Nevertheless, GNN models, while powerful, struggle to capture specific properties like triangles or cycles, thus affecting the performance of our model. Future work will focus on exploring the model’s applicability in various real-world scenarios.

### v3

@@ Abstract
Graph generation has emerged as a crucial task in machine learning, with significant challenges in generating graphs that accurately reflect specific properties. Existing methods often fall short in efficiently addressing this need as they struggle with the high-dimensional complexity and varied nature of graph properties. In this paper, we introduce the Neural Graph Generator (NGG), a novel approach which utilizes conditioned latent diffusion models for graph generation. NGG demonstrates a remarkable capacity to model complex graph patterns, offering control over the graph generation process. NGG employs a variational graph autoencoder for graph compression and a diffusion process in the latent vector space, guided by vectors summarizing graph statistics. We demonstrate NGG’s versatility across various graph generation tasks, showing its capability to capture desired graph properties and generalize to unseen graphs. We also compare our generator to the graph generation capabilities of different LLMs. This work signifies a shift in graph generation methodologies, offering a more practical and efficient solution for generating diverse graphs with specific characteristics.
@@ 1 Introduction
Recent graph generative models have improved in accurately representing various properties of real-world graphs. Nonetheless, to produce graphs adhering to specific properties, it is necessary first to compile a dataset of graphs that display these characteristics before proceeding to train the model using this dataset. Thus, to generate different types of graphs, different models need to be trained, which is impractical. This work aims to fill this gap. Specifically, we develop a neural network model, so-called Neural Graph Generator (NGG), to generate graphs that exhibit specific properties. We capitalize on recent advances in latent diffusion models to perform vector-conditioned graph generation. The model applies the diffusion process not to the graph data but instead to an encoded latent representation of the graph. We produce vectors that contain a summary of the statistics of each graph (e. g., number of nodes, number of edges, clustering coefficient, etc.). Those vectors guide the denoising part of the diffusion process. We first pre-train a variational graph autoencoder which we use to map graphs into vectors and map vectors into graphs. The diffusion is performed in the space of vectors. To generate new graphs that exhibit specific properties, we sample a vector from the standard normal distribution, and we denoise this vector, while denoising is guided by the vector of graph properties. Once the noise has been removed from the vector, we feed the emerging vector into the decoder of the pre-trained variational graph autoencoder to produce a graph. The proposed model is extensively evaluated on two axes: i. tasks of graph generation conditioned on multiple graph properties ii. comparison to graph generation results of different state-of-the-art LLMs (Llama2, Llama3, Gemma, Mistral and GPT3.5). Our results indicate that the NGG pretrained model can indeed generate graphs whose properties are similar to the ones that are given as input to the model and also it out performs LLMs in all aspects including error, accuracy and execution times where our approach is orders of magnitude faster than LLMs while it consumes much less memory in GPU.
To summarize, our work makes the following contributions:
@@
    we perform - as part of the experimental evaluation of NGG - the first comprehensive evaluation of LLMs in graph generation tasks, covering many generic properties of graphs.
  • (v)
@@ 2 Related Work
For the problem of graph generation, as discussed above, most models belong to one of the following five families: (1) Auto-regressive models; (2) Variational autoencoders; (3) Generative adversarial networks; (4) Normalizing flows; and (5) Diffusion models. Auto-regressive models assume a specific node ordering and generate graphs in a sequential fashion. GraphRNN [You et al., 2018], GraphGen [Goyal et al., 2020] and GraphGen-Redux [Bacciu and Podda, 2021] are three examples of auto-regressive models. GraphRNN generates nodes and its associated edges sequentially, while the other two models generate edges and their endpoint nodes sequentially. Variational autoencoders consist of two modules. First, the encoder which maps the input data to a space that corresponds to the parameters of a Gaussian distribution. Typically, this distribution is encouraged to be similar to the standard normal distribution. Second, the decoder which maps from the latent space to the input space. In the graph domain, GraphVAE [Simonovsky and Komodakis, 2018] is a model whose encoder is an instance of a message passing graph neural network, while its decoder is a simple multi-layer perceptron. There are models that embed graphs into distributions different from Gaussian such as DGVAE [Li et al., 2020] which utilizes the Dirichlet distributions as priors on the latent variables and the latent variables represent graph cluster memberships. Variational autoencoders have also achieved success in the field of chemoinformatics [Jin et al., 2018, Samanta et al., 2020]. Similar to variational autoencoders, generative adversarial networks also consist of two main components, a generator whose objective is to generate realistic samples and a discriminator whose objective is to distinguish real samples from synthetic ones. Most models from this family typically consist of a generator that takes vectors sampled from a standard normal distribution and processes them with a multi-layer perceptron to generate a graph, and of a discriminator which is an instance of a message passing graph neural network [De Cao and Kipf, 2018, Pölsterl and Wachinger, 2021]. Architectures based on normalizing flows explicitly model a probability distribution by leveraging a method that uses the change-of-variable law of probabilities to transform a simple distribution into a complex one by a sequence of invertible and differentiable mappings. Normalizing flows can be applied to graphs by designing message passing mechanisms that are exactly reversible [Liu et al., 2019]. Such models have been applied to molecular data. For example, MoFlow [Zang and Wang, 2020] generates chemical bonds through a Glow-based model, and atoms given bonds using a graph conditional flow. Diffusion models gradually add noise to the input data in the forward diffusion process, and then learn to remove the noise in the reverse diffusion process. There exist models that add Gaussian noise to the graph’s adjacency matrix and binarize the continuous values of the output of the reverse diffusion process to produce valid graphs [Niu et al., 2020, Jo et al., 2022]. More recent models apply a discrete diffusion process that progressively adds noise by adding or removing edges and changing the nodes’ and/or edges’ types [Vignac et al., 2023]. Recently, several studies have focused on the potential of large language models (LLMs) for graph generation, demonstrating promising preliminary capabilities in both rule-based and distribution-based graph generation [Yao et al., 2024].
@@ 3 Neural Graph Generator
Figure 1: Overview of the proposed architecture. The variational graph autoencoder is responsible for generating a compressed latent representation $\boldsymbol{z}$ for each graph $\boldsymbol{G}$. Those representations are fed to the diffusion model which operates in that latent space adding noise to $\boldsymbol{z}$ resulting to $\boldsymbol{z}_{T}$ The denoising process is conditioned on the encoding (output of $\tau_{\theta}$) of the vector that contains the graph’s properties. The output of the diffusion model is passed on to the decoder which generates a graph.
@@ 4 Experimental Evaluation
With regard to the variational autoencoder, the encoder $\mathcal{E}$ consists of $2$ layers, the hidden dimension size is set equal to $64$, while the input graphs are embedded into $32$-dimensional vectors. We imposed an ordering on the nodes based on the BFS tree’s default ordering. The decoder $\mathcal{D}$ consists of $3$ layers and the hidden dimension size is set to $256$. The output of $\mathcal{D}$ is a $100^{2}$-dimensional vector and is transformed into an adjacency matrix of dimension $(100\times 100)$. Thus, the model can handle graphs of up to $100$ nodes. We also initially annotate the nodes of each graph with $10$-dimensional spectral features (from the eigenvectors associated with the $10$ smallest eigenvalues of the normalized Laplacian matrix). For the latent diffusion model, we set the number of timesteps $T$ equal to $500$, while we use a cosine schedule for adding noise at each time step. The denoising neural network $\epsilon_{\theta}$ consists of $3$ layers, the hidden dimension size is equal to $512$, while the size of the output is set to $128$. To train both models, we use the Adam optimizer and we set the learning rate to $0.001$. The batch size is set to $256$. We set the number of epochs of the variational autoencoder to $200$, while the number of epochs of the latent diffusion model to $100$. Overall, the variational autoencoder consists of $2,640,492$ parameters, while the diffusion model consists of $973,088$ parameters. Thus, in total, the NGG model consists of $3,613,580$ parameters. The model was pretrained for about 24 hours in a GPU RTX A6000.
@@
We also use a variant of the variational autoencoder (VGAE) as a baseline where the process is conditioned on vector of properties $\mathbf{c}$ for controlled sampling by concatenating the vector with the graph’s latent representation. Additionally, we conducted further experiments on a subset of the original test set, comparing the performance of the variational autoencoder (VGAE) and our model against a range of LLMs.
@@
Assortativity coefficient 0.30 59.14 0.11 37.71
# triangles 9,284.64 89.57 920.24 21.86
Avg. # triangles formed by an edge 7.94 59.64 8.46 48.46
Max. # triangles formed by an edge 483.26 75.28 44.17 14.94
Avg. local clustering coefficient 0.27 31.42 0.07 13.82
Global clustering coefficient 0.32 52.10 0.05 12.61
@@ 4.1 Within Distribution Performance
# nodes 15.49 18.77 6.69 6.67 17.89 18.13 19.28 19.45
# edges 355.79 60.13 99.09 11.32 548.98 34.84 585.83 37.24
Density 0.29 48.10 0.07 10.74 0.19 23.63 0.20 25.68
Min. degree 12.59 67.12 7.95 41.49 9.70 40.09 9.31 40.07
Max degree 4.46 7.95 2.94 4.64 16.51 24.22 18.14 26.52
Avg. degree 14.35 51.24 3.12 9.27 13.17 28.56 14.22 30.86
Assortativity coefficient 0.29 63.52 0.70 46.01 0.31 40.69 0.29 43.95
# triangles 3,562.78 76.28 1,349.25 24.77 9,369.57 46.48 9,941.35 49.80
Avg . # triangles formed by an edge 7.08 50.06 8.39 50.41 11.18 46.72 11.13 48.45
Max . # triangles formed by an edge 253.39 58.72 75.69 16.12 437.86 66.55 477.43 44.61
Avg. local clustering coefficient 0.24 23.68 0.08 14.48 0.21 25.43 0.23 27.73
Global clustering coefficient 0.25 42.03 0.06 13.15 0.20 25.98 0.21 27.72
Max k-core 11.61 52.76 2.28 9.99 11.79 42.96 17.99 32.05
# communities 2.65 27.47 1.02 12.32 1.29 15.82 1.36 16.24
Diameter 3.53 29.99 2.55 16.62 2.35 16.44 2.31 16.89
All 1.76 56.22 1.23 29.01 0.77 42.16 0.78 42.91
@@ 4.3 Conditioning on Subset of Properties
4.4 NGG vs LLMs
Recent work has highlighted the expressiveness of large language models (LLMs) [Chen et al., 2024], particularly in tasks related to graph generation. In this study, we compared our model and a conditional VGAE against seven LLMs (Llama2-7B, Llama2-13B, Llama-8B, Gemma1.1-2B, Gemma1.1-7B, Mistral-7B, and GPT-3.5-Instruct), evaluating their performance in graph generation, execution time, and GPU usage. For the experiments, we provided each LLM with the required graph properties and prompted it with generating an edgelist that adhered to these specifications. We conducted both zero-shot and one-shot experiments. In the one-shot scenario, we presented the LLM with the edgelist of a randomly chosen graph with fewer than 20 nodes, along with its 15 properties (same as the ones in the condition vector used above), as an example. Tables 3 and 4 present the performance of the LLMs, conditional VGAE, and our model (NGG) on all graph properties. The evaluation was conducted on a subset of 1,000 samples, drawn from the original test set of 100K graphs, due to time and resource constraints. The subset was selected such that it reflects the full dataset graph types distribution. The results in the tables clearly demonstrate that both VGAE and NGG outperform the LLMs across all evaluated criteria (error, execution time and GPU memory usage), with NGG emerging as the top performer. This illustrated the advantages of the NGG model, achieving superior generative performance but also operating more efficiently, as it is about $10^{3}$ times faster than LLMs and at the same time consuming fewer resources and being more environmentally friendly. Additionally, it is noteworthy that the one-shot approach underperformed compared to zero-shot for all Llama models, while for Gemma1.1, Mistral, and GPT-3.5, the one-shot method improved the quality of the generated edgelists.
Model MAE SMAPE (%) Execution Time GPU usage(MB)
Llama2-13B 3.49 56.01 12h & 27min 8,883
Llama2-7B 2.98 54.21 14h & 17min 4,943
Llama3-8B 3.32 50.51 1d & 10h 6,376
Gemma1.1-2B 2.60 57.59 03h & 38min 2,455
Gemma1.1-7B 2.96 57.64 02h & 42min 6,389
Mistral-7B 4.35 58.14 08h & 51min 4,896
GPT-3.5-Instruct 7.18 52.50 04h & 40min -
VGAE 1.44 48.97 16s 720
NGG (ours) 1.31 43.78 16s 800
Table 3: Model performance across all graph properties, evaluated on a subset of 1,000 samples from the initial test set (LLMs evaluated with zero-shot prompts).
Model MAE SMAPE (%) Execution Time GPU usage(MB)
Llama2-13B 12.35 57.28 10h & 46min 6,722
Llama2-7B 12.34 57.09 8h & 33min 5,617
Llama3-8B 5.74 51.16 3d & 20h 7,818
Gemma1.1-2B 2.49 59.00 1h & 04min 2,830
Gemma1.1-7B 2.87 55.39 4h & 57min 9,242
Mistral-7B 2.39 47.55 1d & 20h 6,873
GPT-3.5-Instruct 2.69 57.03 01h & 44min -
VGAE 1.44 48.97 16s 720
NGG (ours) 1.31 43.78 16s 800
Table 4: Model performance across all graph properties, evaluated on a subset of 1,000 samples from the initial test set (LLMs evaluated with one-shot prompts).
@@ 4.6 Uniqueness of Generated Graphs
4.7 Limitation of baselines
In our work, we chose not to conduct experiments comparing our model with the DiGress [Vignac et al., 2023] or EDGE [Chen et al., 2023] or other state-of-the-art generative models. The primary reason for this decision is that these architectures are not compatible with the conditioning/prompting mechanism we propose involving multiple fuatures of a graph.
If we would want to compare directly we should adapt these models to our specific conditioning mechanism and synthetic dataset. Implementing such adaptations would not only be technically demanding but would also represent a new contribution, beyond the scope of our current study. We leave it as potential future work though that could explore these adaptations in detail, potentially leading to novel contributions in the application of DiGress and EDGE to similar datasets.
@@ 5 Conclusion
We introduced the Neural Graph Generator (NGG), a conditional latent diffusion model, for efficient and accurate graph generation. NGG represents a significant new paradigm, adeptly generating graphs conditioned on a vector of multiple specific, user-defined properties, a task that has long challenged traditional models. We demonstrated the ability of NGG to generalize to graphs beyond those in the training set, and the ability to handle missing values in the condition vector. We also highlighted the importance of a dedicated graph generator, particularly in scenarios where LLMs struggle to capture complex structural representations. Our model not only outperforms LLMs in graph generation tasks but also does so more efficiently, consuming fewer resources. Nevertheless, GNN models, while powerful, struggle to capture specific properties like triangles or cycles, thus affecting the performance of our model. Future work will focus on exploring the NGG’s extensions/adaptations to various real-world scenarios (i. e., power/telecom networks, molecules/protein graphs), and scaling it to larger graphs.
@@ Appendix A Dataset Details
Table 6: Description of the $15$ considered graph properties.
Appendix B LLM prompting examples
In tables 3 and 4, it is clear that we employed different output lengths in terms of tokens, varying both by the specific LLM used and by the learning approach (zero-shot or one-shot). This variation was essential to optimize performance within the constraints of each model and task. The smaller LLMs require less GPU memory, allowing us to generate more tokens during inference. For models like ChatGPT, we determined that the output length is governed by the formula: output length = 4k - input length. Consequently, the maximum allowable output length in each scenario is constrained by the available resources or API limitations.
Figures 3(a) and 3(b) provide examples of generalized prompting for zero-shot and one-shot learning, respectively.
(a) Example of zero-shot prompting
(b) Example of one-shot prompting
Figure 3: Prompting examples for LLM graph generation
Appendix C Ablation Study
C.1 Different methods for the construction of adjacency
Table 7 illustrates the impact of different node ordering techniques on the construction of the adjacency matrix and the resulting performance of the model. The results indicate that initiating the BFS algorithm from the node with the highest degree is the most effective approach. This method consistently leads to better performance, highlighting its prominence in structuring the adjacency matrix. Interestingly, when the adjacency matrix is constructed starting from the node with the highest PageRank score, there is a risk of generating a few invalid graphs during inference (about $0.002\%$ invalid graphs). This suggests that while PageRank-based ordering can be useful, it may introduce complications that undermine the model’s reliability. Additionally, ordering the nodes based on degree without the BFS algorithm also yields good results, but it slightly underperforms compared to the BFS-based approach. This observation underscores the advantage of using BFS with degree-based root selection in optimizing graph representation and model accuracy.
Method MAE SMAPE (%) Percentage of Non-Valid Graphs (%)
BFS & Degree 1.05 21.21 0
Degree only 1.78 21.64 0
Pagerank 1.69 22.18 0.002
Table 7: Performance of different ordering methods for the construction of adjacency matrix
C.2 Importance of Graph Properties
Figure 4: Importance of graph properties, given the number of nodes, edges, and one of the remaining 13 properties as input.
Given that we already know the number of nodes and edges, we investigated the significance of individual graph properties when only one of the remaining 13 properties is known (including the number of nodes and edges). In these experiments, we provided the model with the number of nodes, the number of edges, and one additional property, effectively creating a triplet of properties while masking the other 12. We conducted these experiments across the entire test set, repeating each trial three times to minimize the impact of random variation on the results. The mean SMAPE metric from these three runs is reported as the outcome of this ablation study. In Figure 4, the property highlighted in orange is the one that contributes most significantly to the model’s performance. This analysis helps identify which properties, when combined with the number of nodes and edges, have the greatest impact on the model’s performance.